The length of a rectangle is 2 less than 3 times the width. If the perimeter is 68, what is the length of the rectangle?
Answer by acalgebra(30)

------
3w-2=l

perimeter=2(l+w)
2([3w-2]+w)
2(3w-2+w)
2(4w-2)
8w-4

perimeter=68
8w-4=68
8w=72
w=9

3w-2=l
3(9)-2=l
27-2=l
25=l
l=25

So the length is 25.